OpenStudy (anonymous):

If g(x) = x2 − 2, find g(6). Choose one answer. a. 36 b. 10 c. 34 d. −2

OpenStudy (anonymous):

Plug 6 in for x then solve the equation. (6)2-2=g(6)

OpenStudy (anonymous):

D?

OpenStudy (agent0smith):

Is that a guess? Show work.

OpenStudy (anonymous):

They said plug in 6 for x and thats what i did and i got d :P

OpenStudy (anonymous):

at least i think i got d.

OpenStudy (anonymous):

Your supposed to solve the equation. Multiply the first set then subtract 2.

OpenStudy (anonymous):

ok... im so confused though

OpenStudy (agent0smith):

Show us HOW you got -2 from (6)^2 - 2

OpenStudy (anonymous):

g(x) = x^2 - 2 g(6) = 6^2 - 2 = 36 - 2 = 34

OpenStudy (agent0smith):

Good. Notice it's not remotely close to -2 :P

OpenStudy (anonymous):

its c